OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN YOU The volume of a sphere is 1674 m. What is the surface area of the sphere to the nearest tenth?

OpenStudy (anonymous):

\[V_{sphere} = \frac{4}{3}\pi{r^3}\]

OpenStudy (anonymous):

Solve for \(r\) and then plug that into the equation for surface area.

OpenStudy (anonymous):

Sorry im probably being so bothersome right now with my questions. I have a test tomorrow and it is also the end of the grading period I have a B in the class which messes up my straight A status :(

OpenStudy (anonymous):

\[SA_{sphere} = 4\pi{r^2}\]

OpenStudy (anonymous):

Im already confused how do you solve for R

OpenStudy (anonymous):

\[1674 = \frac{4}{3}\pi{r^3}\]

OpenStudy (anonymous):

divide 1674 by 4/3?

OpenStudy (anonymous):

Yes. And then divide that by \(\pi\), and take the cube root.

OpenStudy (anonymous):

I got 7.365

OpenStudy (anonymous):

That's the correct value for \(r\).

OpenStudy (anonymous):

Alright so 4xpix7.3

OpenStudy (anonymous):

Don't forget that it's \(r^2\).

OpenStudy (anonymous):

670

OpenStudy (anonymous):

rounded

OpenStudy (anonymous):

You lost some accuracy from just using 7.3 instead of 7.365 but your answer is right.

OpenStudy (anonymous):

not according to my study guide it should be 1462.5

OpenStudy (anonymous):

Your study guide is wrong: http://www.wolframalpha.com/input/?i=sphere+with+volume+1674

OpenStudy (anonymous):

the study guide is automatically made it cant be wrong..

OpenStudy (anonymous):

But then again my teacher is lazy...

OpenStudy (anonymous):

WolframAlpha isn't wrong.

OpenStudy (anonymous):

If doing it by hand and wolfram alpha both disagree with your study guide it's either a misprint in the question or in the answer.

OpenStudy (anonymous):

Alright thanks anyways. I got like 10 more to go :( and its 1 am looks like no sleep for me lol